OpenStudy (imtiaz7):

The polynomial of x^3+lx^2+mx+24 has a factor (x+4) it leaves the remainder of 36 when divided by (x-2). find the values of l and m.

OpenStudy (anonymous):

hey I solved in this fashion : 1st I divided the given cubic equation with x-2 and solve using simple division and equated its remainder32+2m+4l=36(the given one) and got m+2l=2 and name it equation1. 2nd then I divided the given equation with x+4 and solve using same method and got this equation -40-4m+16l=0(as it is the factor the equation) and solving I get 4l-m=0 and name it equation2.Now solving both equations I get l=2 and m=-2. is it right? @Imtiaz7

OpenStudy (imtiaz7):

no this is not right